OpenStudy (juan1857):

The dimensions in inches of a shipping box at UHAUL can be expressed as width x, length x+5, and height 3x-1. The volume is about 7.6 ft^3. Find the dimensions of the box in inches. Round to the nearest inch.

OpenStudy (math&ing001):

The volume of the box is the product of the width, length, and height. Can you write it in terms of x ?

OpenStudy (juan1857):

??

OpenStudy (math&ing001):

I'll right it in math terms so you can understand better Volume = width * length * height Volume = 7.6 Therefor: width * length * height = 7.6 Replace the width, length, and height with their respective values to get an equation that you can solve for x.

OpenStudy (math&ing001):

write*

OpenStudy (juan1857):

ok, I got 3x^2+14x-5 so how do I divide this from 7.6ft^3?

OpenStudy (math&ing001):

Are sure about 3x^2+14x-5 ? Should be 3-dimentional..

OpenStudy (math&ing001):

Volume = x(x+5)(3x-1) So x(x+5)(3x-1) = 7.6 Distribute and then solve for x.

OpenStudy (juan1857):

oh yeah i didn't multiply the x, do you distribute the x to (x+5) only and then multiply with (3x-1)

OpenStudy (math&ing001):

Yeah whichever way you choose will get you the same thing ;]

OpenStudy (juan1857):

so would it be 3x^3+4x^2-5x?

OpenStudy (juan1857):

oh yeah my bad

OpenStudy (math&ing001):

We have 3x^3 + 14x^2 - 5x = 7.6 So 3x^3 + 14x^2 - 5x - 7.6 = 0 Now we just solve for x

OpenStudy (juan1857):

how do I do that?

OpenStudy (math&ing001):

We need to find a way to factor it..

OpenStudy (juan1857):

sorry, I don't know how, lol

OpenStudy (math&ing001):

I put it on wolfram and this is what I got https://www.wolframalpha.com/input/?i=solve++3x^3+%2B+14x^2+-+5x+-+7.6+%3D+0 Are you sure about the data ?

OpenStudy (math&ing001):

Normally it should be easier to factorize.

OpenStudy (math&ing001):

Oh my bad !!! I forgot to convert to inches !!

OpenStudy (math&ing001):

7.6 ft^3 = ? in^3

OpenStudy (juan1857):

I got lost with the factoring

OpenStudy (math&ing001):

Don't worry about it. 7.6 ft^3 = 13132.8 in^3 So the equation would be 3x^3 + 14x^2 - 5x - 13132.8 = 0 Even here we can't solve it manually... But using a computer we get x=14.972. Meaning: width = x = 14.972 inches length = x+5 = 19.972 inches ...

OpenStudy (juan1857):

ok thanks